I do not know how to plug into calculator
. Evaluate the logarithmic equation for three values of x that are greater than -1, three values of x that are between -2 and -1, and at x = -1. Show your work. Use the resulting ordered pairs to plot the graph; submit the graph via the Dropbox. State the equation of the line asymptotic to the graph.
y = log4 (x + 2)
2. Evaluate the exponential equation for three positive values of x, three negative values of x, and at x=0. Show your work. Use the resulting ordered pairs to plot the graph; submit the graph via the Dropbox. State the equation of the line asymptotic to the graph.
y=3^x-4+1

Answer by jsmallt9(3759)

No calculators I know of can do base 4 logarithms so don't feel bad that you couldn't figure how to use your calculator. The base 4 logarithms can be calculated "by hand" if the argument, x+2, turns out to be a power of 4. So we just have to pick x's so that x+2 works out to be a power of 4.

What are some powers of 4 and how do we figure out an x value that makes x+2 equal to those powers of 4?
  1. Pick an integer (... -3, -2, -1, 0, 1, 2, 3, ...)
  2. Raise 4 to that power.
  3. Set that power of 4 equal to x+2.
  4. Solve for x (i.e. subtract 2 from each side).
  5. Keep track of which x goes with which integer exponent. The integer value will be the base 4 logarithm.
  6. Repeat this process until you have all the x's you need.

Here's some examples:
n     4 to the n power     Equation to find x       x      y = log4(x+2)
                                                            (same as n)
0          1                  x+2 = 1               -1          0
1          4                  x+2 = 4               2           1
-1         1/4                x+2 = 1/4            -1.25       -1
2          16                 x+2 = 16              14          2
-2         1/16               x+2 = 1/6            -1.9375     -2

You need one more x which works out between -1 and -2. I'll leave it up to you to find it.

With these pairs of x and y values (the last two columns) you have ordered pairs which you can plot on a graph. The "connect the dots" with as smooth a curve as possible.

As for the asymptote, logarithmic equations have vertical asymptotes for x values that make the argument of any logarithm zero, if any. To find your asymptote set you argument to zero and solve:
x + 2 = 0
x = -2
So your graph will have a vertical asymptote at x = -2.